OpenStudy (anonymous):

Can you please help me solve this? x-9= 24

OpenStudy (nubeer):

x - 9 + 9 = 24 +9.. can u do this now>?

OpenStudy (anonymous):

First of all you should collect the numbers together and the variables together x=24+9

OpenStudy (anonymous):

x=24+9 =33